If the angle of a projectile is 60 degree, then what is the height of it when the horizontal distance travelled is 3R/4 where R is the range of the projectile


 

Grade:12

1 Answers

Prashant Kumar
6 years ago
Hlo frnds Use equation of trajectoryY= xtanθ(1-x/R)(x = horizontal did.)Substituting the value we get;H=3R/4 * tan60°(1-¾R/R)= 3√3R/4* (1-3/4)=3√3R/16Height = 3√3R/16